A block starts up at point A of an incline of 250 with initial speed v and stops at point B. Calculate speed v if the coefficient of kinetic friction between the incline and the block is µK = 0.40 and d = AB = 2 m.


1
Expert's answer
2021-07-02T09:39:47-0400
"0.5v^2=gl\sin{25}-\mu gl\cos{25}\\
0.5v^2=(9.8)(2)(\sin{25}-0.4 \cos{25})\\v=1.5\frac{m}{s}"